OpenStudy (calculusxy):

Use the distributive property to write the expression as an equivalent algebraic expression. -20(19 - a) Is the answer -180 + 20a?

OpenStudy (alexandervonhumboldt2):

-a*(b-c)=-ab+ac

OpenStudy (calculusxy):

So am I correct?

OpenStudy (calculusxy):

@phi

OpenStudy (alexandervonhumboldt2):

no -20*19=-380

OpenStudy (alexandervonhumboldt2):

-380+20

OpenStudy (alexandervonhumboldt2):

-380+20a sorry